Show calculations for each step.

0.080 mg of Mg reacts with HCl (aq) to give 85.5 mL of og H2 gas at 22oC and 759 mm Hg pressure.

Mg(s) + 2HCl (aq) LaTeX: \longrightarrow

MgCl2 (aq) + H2(g)

a) What is the partial pressure of H2 gas?

b) Use CGL to calculate volume of H2 gas at STP.

(hint: make a table with correct units for lab conditions and STP conditions and then solve for V2).

c) Calculate moles of Mg and get moles of H2 gas using mol-mol ratio from the balanced equation.

d) Calculate molar volume of H2 gas (Hint: divide V2 by moles of H2).

e) Calculate % error (theoretical value = 22.4 L/mol)
1
Expert's answer
2020-05-08T14:11:37-0400

a)

From step (c): n (H2) = (Mg) = 0.00329 mol

PV = nRT

P1 = 759 mm Hg = 101,191.678 Pa

T1 = 22 + 273.15 = 295.15 K

V = nRT / P

V = 0.00329 * 8.314 * 293.15 / 101,191.678

V = 79.78 mL - vlume of H2 in the mixture

79.78 / 85.5 * 101,191.678 Pa = 94.4 kPa - partial pressure of H2

b)

STP: 273.15 K and 101,325 Pa

CGL: P1V1/T1 = P2V2/T2

V2 = P1V1T2/P2V2

P1 = 759 mm Hg = 101,191.678 Pa

V2 = (101,191.678*79.78*(22+273.15))/(101,325*273.15)

V2 = 86.1 mL

c)

M (Mg) = 24.31 g/mol

n = 0.080 / 24.31 = 0.00329 mol

d)

V2 = 86.1 mL

Vm = 86.1 mL / 0.00329 mol = 26.2 (L/mol)

e)

Theoretical value = 22.4 L/mol

% error = (26.2 - 22.4)/22.4 = 0.17 or 17%